OpenStudy (anonymous):

integral of (cos theta)^2 d theta?

OpenStudy (anonymous):

\[\int\limits \text{Cos}[\theta ]^2 \, d\theta \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

could you please look at my previous question? I type the notation there

OpenStudy (anonymous):

sorry imrammeah, it was wrong

OpenStudy (anonymous):

trick is to write \[\cos^2(\theta)=\frac{1}{2} \cos(2\theta)+\frac{1}{2}\]

OpenStudy (anonymous):

could you please explain, where did you derive that from?

OpenStudy (anonymous):

back of a text book

OpenStudy (anonymous):

that is how i do all integration problems because they are stultifyingly boring tricks

OpenStudy (anonymous):

\[\color{blue}{\text{myininaya will explain}}\]

myininaya (myininaya):

ok do you remember that cos(x+x)=cos(2x)=cosxcosx-sinxsinx=cos^2x-sin^2x (but sin^2x=1-cos^2x) cos^2x-sin^2x=cos^2x-(1-cos^2x)=2cos^2x+1 ok so we have cos(2x)=2cos^2x+1 solve this for cos^2x

OpenStudy (anonymous):

see! you think i can remember that?

myininaya (myininaya):

ok and i made a little type up there cos2x=2cos^2x-1

OpenStudy (anonymous):

use U substitution. \[\int\limits_{}^{} (u)^{2} du\]\[\cos \theta=u\]\[du=-\sin \theta\] \[=u^3/3\]\[=(-\sin \theta)/3\]

OpenStudy (anonymous):

rsmith, where do I find sin theta in the problem?

OpenStudy (anonymous):

guys, why can't I find the "good answer" button anywhere?

OpenStudy (anonymous):

using U substitution you take the derivative of what you take out for U hence the cos theta becoming -sin theta

myininaya (myininaya):

\[\int\limits_{}^{} (\frac{1}{2}\cos(2x)+\frac{1}{2}) dx=\frac{1}{4}\sin(2x)+\frac{1}{2}x+C\]

OpenStudy (anonymous):

ok, is this the answer:\[1/2\theta + 1/4 \sin 2\theta + C\]

myininaya (myininaya):

yes!

OpenStudy (anonymous):

ok, thank you, still cant find the button

OpenStudy (anonymous):

well then. haha i guess i was wrong :(

myininaya (myininaya):

^^^^^ |||||||| the button is here

OpenStudy (anonymous):

i'll give her one

OpenStudy (anonymous):

btw if you are doing lots of these the back of your text is really your best friend. other than here of course

OpenStudy (anonymous):

thanks for the tip.

OpenStudy (anonymous):

hey aftercall, does your book have the answers in them?